OpenStudy (anonymous):

Determine whether there exists an integrating factor that is a function of x only that would make the following differential equations exact. If so, compute the integrating factor and determine the solution to the differential equation. (7x^2)*(sqrt(x*y))+cos(x*y)dy/dx=0 y(0)=0

OpenStudy (anonymous):

I don't think there is an integrating factor that is a function of x only for that differential equation. In my DE book I found a section which says if \[(\delta P/ \delta y - \delta Q/ \delta x)/Q\] is a function of x only, then \[e ^{\int\limits_{}^{}h(x)dx}\] is an integrating factor where \[Pdx + Qdy = 0\] is the form of the differential equation. I calculated this and did not get a function of only x. Check my math but that seems to answer your question.

OpenStudy (unklerhaukus):

\[\begin{align} \text{Integrating factor } R =R(x) \\ \\ \qquad&\text{for exactness } \\ \\ \frac{\partial}{\partial y}\big(R(x)P\big) &=\frac{\partial}{\partial x}\big(R(x)Q\big) \\ RP_y &=R'Q+RQ_y \\ R(P_y-Q_x) &=R'Q \\ \frac {R'}R &=\frac{P_y-Q_x}Q \\ \int\frac{d R}R&=\int\frac{P_y-Q_x}Qdx \\ \ln Q &=\int\frac{P_y-Q_x}Qd x \\ R(x) &=\Large e^ {\int\frac{P_y-Q_x}Qd x} \end{align}\] But i think your right @JonathanHocker, this only works iff \[\frac{P_y-Q_x}Q=h(x) \] a function of x only. but that dosen't seam to be the case here